OpenStudy (anonymous):

Given the sample triangle below and the conditions : sin B =1/2, A= 20. find the hypotenuse of the triangle A. SqaureRoot3/2 B. 40SqaureRoot3/2 C.2 D.40

OpenStudy (triciaal):

try to do like the one I walked you through earlier . remember it's not hard.

OpenStudy (anonymous):

This is a right triangle, so we can use the definition of sin from SOH CAH TOA: Sin (of angle) = Opposite side / hypotenuse Sin of angle B = 1/2 (from problem statement) Sin of angle B = opposite / hypotenuse = b/c We know therefore that b/c = 1/2, and we can use that with a = 20 and the Pythagorean theorem to find b and c.

OpenStudy (anonymous):

Solve bc=12 for b in terms of c. Then write out a2+b2=c2 substitute in your value for b in terms of c on the left hand side, and the value of a=20 on the left hand side. You should then have an equation in terms only of c to solve, and the positive answer will be the length of the hypotenuse.

OpenStudy (anonymous):

sorry its a^2 + b^2 = c^2

OpenStudy (anonymous):

yea i know. but where did 12 come from?

OpenStudy (anonymous):

SORRY ITS \[B/C = 1/2\]

OpenStudy (anonymous):

ok then whats next?

OpenStudy (triciaal):

set up your triangle the hypotenuse is opposite the right angle use pythagorean theorem c^2 = a^2 + b^2 sin is O/H

OpenStudy (anonymous):

i have mentioned it in an above reply

OpenStudy (triciaal):

@ChiefArnav the computer has been messing up. I walked her through this earlier also.

OpenStudy (anonymous):

hmmm then i may leave it up to you?

OpenStudy (anonymous):

but i think i walked her up the whole problem.

OpenStudy (anonymous):

I get the A^2 + b^2 = c^2 thing but I dont get how I put b/c = 1/2 in that problem

OpenStudy (triciaal):

@ChiefArnav no please finish, stay

OpenStudy (triciaal):